OpenStudy (anonymous):

Suppose that the amount of air in a balloon after t hours is given by V(t)=t3 -6t2 +35 Estimate the instantaneous rate of change of the volume after 5 hours. Solution Find the ARC first: A.R.C.=V(t)-V(5)/ t-5, so t^3 -6t^2 +35-10/ t-5, so, t^3-6t^2+25 How did he get 10?

OpenStudy (anonymous):

I cant attach a file instead

OpenStudy (anonymous):

hope that makes it easier to read

myininaya (myininaya):

\[\lim_{t \rightarrow 5}\frac{V(t)-V(5)}{t-5}\] \[=\lim_{t \rightarrow 5}\frac{(t^3-6t^2+35)-(5^3-6(5)^2+35)}{t-5}\] Now you have like terms on top simplify that for me and i will be back

OpenStudy (anonymous):

ok that works

OpenStudy (anonymous):

thanks

myininaya (myininaya):

oh you just didn't understand how to plug in?

OpenStudy (anonymous):

yeah pretty much; got confused from another exxample

OpenStudy (anonymous):

thanks :$

myininaya (myininaya):

ok and you are gonna want to approximate the instantaneous rate of change (above if you know limits you can find it exactly) So just take values really close to 5 for t to find an approximation of the i.r.c.

OpenStudy (anonymous):

yup learning limits atm